OpenStudy (anonymous):

choose the ordered pair (x,y) that is a solution for the equation -7x - 14y = -98

OpenStudy (mathteacher1729):

This is easy. Let x = 0 and solve for y. Or let y = 0 and solve for x. :)

OpenStudy (anonymous):

i haven't learned how to do this yet though

OpenStudy (mathteacher1729):

It's not too hard. -7x - 14y = -98 -7(0) - 14y = -98 let x = 0 0 -14y = -98 -14y = 98 y = 98/(-14) divide by -14 y = -7 simplify One ordered pair is (0,-7) Can you find the ordered pair if you let y = 0 first?